OpenStudy (anonymous):

What is the solution to the equation 1 over the square root of 8 = 4(m + 3)

OpenStudy (abb0t):

\[\frac{ 1 }{ \sqrt{8} } = 4(m+3)\]

OpenStudy (abb0t):

is that correct?

OpenStudy (anonymous):

m = –15 over 4 m = –11 over 4 m = 5 over 4 m = 9 over 4

OpenStudy (abb0t):

4m+12 = \(\frac{1}{\sqrt{8}}\), subtract by 12, and then divide by 4.

OpenStudy (abb0t):

by the way, \(\sf \color{red}{\sqrt{8}=2\sqrt{2}}\)

OpenStudy (abb0t):

You might want to reduce that, it helps.

OpenStudy (anonymous):

thanks